OpenStudy (anonymous):

A triangle has side lengths of 9 in, 13 in, and 20 in. What is the measurement of this triangle’s largest angle? 29.9° 50.1° 129.9° 136.2°

OpenStudy (wolf1728):

When we know all 3 sides, we use the Law of Cosines: The largest side is 20 so C is the largest angle cos(C) = (a^2 + b^2 -c^2) / (2*a*b) cos (C) = (9^2 + 13^2 -20^2)/ =2*9*13 cos (C) = (81 + 169 -400) / 234 cos (C) = -150 / 234 cos (C) = -0.641025641 The arc cosine of -0.641025641 is 129.87 Degrees

OpenStudy (anonymous):

that is correct